### v1

# Frances Dean    Jenna Fields    Radhika Bhalerao    Marie Charpignon    Ahmed Alaa

Abstract
Instrumental variable (IV) methods mitigate bias from unobserved confounding in observational causal inference but rely on the availability of a valid instrument, which can often be difficult or infeasible to identify in practice. In this paper, we propose a representation learning approach that constructs instrumental representations from observed covariates, which enable IV-based estimation even in the absence of an explicit instrument. Our model (ZNet) achieves this through an architecture that mirrors the structural causal model of IVs; it decomposes the ambient feature space into confounding and instrumental components, and is trained by enforcing empirical moment conditions corresponding to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). Importantly, ZNet is compatible with a wide range of downstream two-stage IV estimators of causal effects. Our experiments demonstrate that ZNet can (i) recover ground-truth instruments when they already exist in the ambient feature space and (ii) construct latent instruments in the embedding space when no explicit IVs are available. This suggests that ZNet can be used as a “plug-and-play” module for causal inference in general observational settings, regardless of whether the (untestable) assumption of unconfoundedness is satisfied.
1 Introduction
Instrumental variable (IV) regression is a commonly used approach to address unobserved confounding when estimating causal effects from observational data. An IV randomizes the treatment without having a direct effect on the outcome, enabling estimation of causal effects with two stage regression (Imbens and Rubin, 2015). For example, in economics, geographical proximity to a college is used as an instrument for educational attainment in estimating the returns of schooling (Card, 1993) and draft lotteries as instruments to study the effect of military service on long-term economic outcomes (Angrist, 1990). In medicine, genetic variants have been used as IVs since they often highly correlate with risk factors but do not directly influence the outcomes associated with these factors (Davey Smith and Hemani, 2014). While IV regression enables causal identification in observational data, these methods are not ubiquitous because a suitable instrument must be available and known to the analyst. This is seldom the case in many applications. For instance, in health research, genetic variants are rarely available in large datasets, and moreover, candidate genes may not be strong or valid IVs (Davies et al., 2015; Burgess et al., 2016).
Figure 1: Causal inference with learned instruments. We learn a feature representation of the observed data $X$ that decomposes them into a learned instrument $\widetilde{Z}=g(X)$ and a residual confounder $\widetilde{X}=f(X)$. This decomposition is induced by enforcing the moment conditions needed for the learned instrument $\widetilde{Z}$ to be valid.
The tradition of IV regression originated in settings with tabular data, where an instrument is expected to be explicitly observed within the feature space. In contrast, the increasing use of deep learning representations of high-dimensional data, such as text, images, and multimodal inputs, creates new opportunities to uncover instrumental variables that are implicit rather than directly observed. For example, medical images may encode subtle provider- or institution-specific patterns, while rich electronic health records may imitate genetic or phenotypic information through descriptions embedded in clinical notes. Representation learning methods could be used to extract such latent variables as instruments, thereby automating instrument selection and further extending IV methods to non-tabular settings where no single variable in the ambient feature space is a valid IV on its own.
In this paper, we develop a representation learning approach that constructs IVs from observed data by decomposing the input feature space into confounding and instrumental components (Fig. 1). Our model, which we call ZNet, is an encoder architecture that mirrors the structural causal model (SCM) of IVs and is regularized by enforcing the empirical moment conditions that correspond to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). The learned instrumental representation can then be used as input to standard two-stage IV estimators to recover causal effects. When a valid IV is present in the ambient feature space, ZNet recovers a representation that is strongly correlated with it; when no explicit instrument exists, ZNet learns a latent representation that can serve as a pragmatic instrument. The ZNet architecture is compatible with a broad class of downstream IV estimators, and through semi-synthetic experiments we show that learning instrumental representations substantially reduces bias from unobserved confounding. Additionally, we illustrate the learned feature decomposition in real-world, high-dimensional settings using electrocardiogram data.
2 Preliminaries
Let $Y\in\mathcal{Y}\subset\mathbb{R}$ denote a continuous outcome, $T\in\mathcal{T}\subset\mathbb{R}$ a treatment variable, and $X\in\mathcal{X}\subset\mathbb{R}^{d}$ a set of observed covariates associated with each unit. The treatment $T$ has a causal effect on the outcome $Y$, while $X$ may influence both $T$ and $Y$ ($X$ confound the relation of primary interest between $T$ and $Y$). In addition to the measured confounders $X$, there are unknown or unobserved confounders $U$, which induce spurious associations by simultaneously affecting both the treatment and outcome (Fig. 1). We assume that the outcome variable $Y$ is determined by the following SCM:
$Y=\varphi(X,T)+e_{Y}(U),\,\,\,\,\,\,\,\,T=\psi(X)+e_{T}(U),$ (1)
for unknown functions $\varphi:\mathcal{X}\times\mathcal{T}\rightarrow\mathcal{Y}$ and $\psi:\mathcal{X}\rightarrow\mathcal{T}$. Following (Hartford et al., 2017), we assume that the unobserved confounders $U$ influence the outcome $Y$ and the treatment $T$ additively, via the error functions of U, which we henceforth denote $e_{Y}$ and $e_{T}$. As a result, the observational and interventional distributions generally differ, i.e.,
$\mathbb{E}[Y|X,T]$ $=\varphi(X,T)+\mathbb{E}[e_{Y}|X,T]$
$\neq\varphi(X,T)+\mathbb{E}[e_{Y}|X]=\mathbb{E}[Y|do(T),X].$
Thus, standard regression would lead to confounding bias.
De-confounding with instrumental variables (IVs). A common method for removing confounding bias is to use IV regression. In the classical IV setting, we assume access to an additional variable $Z$ that is not influenced by the unobserved confounders $U$, affects the treatment $T$, and has no direct effect on the outcome $Y$ (Verbeek, 2004; Angrist et al., 1996). Formally, given a set of observed confounders $X$, $Z$ is a valid IV if it satisfies the following conditions:
Relevance: $\,\,Z\not\perp T\,|\,X,$
Exclusion restriction: $\,\,Z\perp Y\,|\,(X,T,U),$
Unconfoundedness: $\,\,Z\perp U\,|\,X.$ (2)
Under these conditions, the instrument $Z$ can be used in a two-stage regression to estimate the effect of $T$ on $Y$. Under the additive model in (1), Hartford et al. (2017) uses the instrument $Z$ to set up an inverse problem by relating the counterfactual $\mathbb{E}[Y|do(T),X]$ to observable distributions:
$\mathbb{E}[Y|X,Z]$ $=\mathbb{E}[\varphi(X,T)+e_{Y}|X,Z]$
$=\mathbb{E}[\varphi(X,T)|X,Z]+\mathbb{E}[e_{Y}|X]$
$=\mbox{$\int$}\,\mathbb{E}[Y|do(T),X]\,dF(T|X,Z).$ (3)
Thus, with the instrument $Z$, we can estimate the counterfactual $\mathbb{E}[Y|do(T),X]$ by estimating the two observable functions $\mathbb{E}[Y|X,Z]$ and $F(T|X,Z)$. While this inverse problem is ill-posed, it provides a practical framework for estimating counterfactuals, and identification is possible under certain conditions (Newey and Powell, 2003). [^1] For example, in the linear case, Two-Stage Least Squares Regression (TSLS) allows for identification of causal effects. A two-stage regression fits a model $\widehat{F}(T|X,Z)$, and estimates $\mathbb{E}[Y|do(T),X]$ by replacing $F(T|X,Z)$ with $\widehat{F}(T|X,Z)$ in (2). With an estiimate of $\mathbb{E}[Y|do(T),X]$, we can derive the conditional average treatment effects (CATE) and average treatment effects (ATE):
$\operatorname{CATE}(X)$ $=\mathbb{E}[Y|do(T)=1,X]-\mathbb{E}[Y|do(T)=0,X],$
$\operatorname{ATE}$ $=\mathbb{E}[Y|do(T)=1]-\mathbb{E}[Y|do(T)=0].$ (4)
Equation (2) is notably more general than IV regression with linear models. We allow $\mathbb{E}[e_{Y}|X]\neq 0$, i.e., observed confounders can correlate with unobserved errors. In two stage least square regression (TSLS), confounders cannot be endogenous for unbiased estimates (Verbeek, 2004).
Figure 2: Illustration of the IV setting. (a) Causal graph for the IV setting; (b) When an IV exists, causal effects can be estimated under unobserved confounding by first fitting a model $\psi$ to predict $T$ from $(X,Z)$, and then training a model $\varphi$ that predicts outcomes based on predicted treatments. (c) Illustration for the latent instrument example discussed in Section 3. Here, $P$ is the provider identity.
3 Learning Instruments from Data
In the absence of a (known) valid instrument $Z$, our method learns an instrumental representation directly from the features $X$ using an observational dataset $\{(X_{i},T_{i},Y_{i})\}_{i}$.
Instrument-confounder decompositions. As illustrated in Fig. 1, we seek a pair of functions $(f(\cdot),g(\cdot))$ that decompose the observed features $\widetilde{X}$ into an observed confounding component $C$ and an instrumental component $\widetilde{Z}$, i.e.,
$(\widetilde{X},\widetilde{Z})=(f(X),g(X)),$ (5)
such that $g(X)$ satisfies the three IV conditions in (2), and the function $f(X)$ fits the additive SCM in (1), i.e.,
$Y=\varphi(f(X),T)+e_{Y}(U),\,\,\,\,T=\psi(f(X))+e_{T}(U).$
Notice that this construction is broadly applicable and captures many real-world scenarios for how instrumental variations may be encoded in the feature space. First, consider the scenario where a subset of the observed variables can already serve as a valid instrument, i.e., $Z\subset X$. In this setting, the decomposition trivially reduces to $g(X)=Z$ and $f(X)=X\setminus Z$. Second, consider scenarios in which the instrument is latent but can be inferred from proxy signals in the observed features $X$. For example, in healthcare settings, providers are often quasi-randomly assigned and have different propensities to administer treatments based on their interpretation of patient information, while provider identity itself does not directly affect outcomes. Although provider identifiers may not be explicitly recorded, their influence shows in retrospective data through visit timing, laboratory and diagnostic tests ordered, and clinical notes (Fig. 2(c)). A representation of provider influence can therefore be learned from $X$ via $g(\cdot)$ and used as an instrumental variable.
While the instrument–confounder decomposition is motivated by real-world scenarios, our model does not require an a posteriori interpretation of the learned instrument $\widetilde{Z}$. Satisfaction of the three IV conditions allows for a pragmatic implementation of IV regression independent of interpretation, since our decomposition treats instruments as abstract mathematical constructs in a latent embedding space, even when no explicitly identifiable instrumental variable exists. Consequently, our method can be applied without prior domain knowledge on the existence of valid instruments.
Moment conditions on $f(.)$ and $g(.)$. A valid decomposition of the observed features $X$ into confounding and instrumental components must induce a factorization of the joint distribution $\mathbb{P}(g(X),f(X),U,T,Y)$ that is consistent with the SCM defined by (1) and (2). In particular, the relevance condition requires that the instrumental component $g(X)$ be predictive of the treatment, i.e., $\operatorname{Cov}(g(X),T)\neq 0$. Exclusion restriction requires that all direct effects of $X$ on the outcome $Y$ be mediated through the confounding component $\widetilde{X}=f(X)$; equivalently, $\operatorname{Cov}(f(X),Y)\neq 0$ while $\operatorname{Cov}(f(X),g(X))=0$. As we discuss later in detail in Section 5, we enforce these covariance relationships through the architecture and loss function of our proposed model.
The covariance structure described above ensures that the learned instrument $\widetilde{Z}$ will be unconfounded and independent of the error $e_{Y}$ as long as $X$ is unconfounded by $U$. The assumption that $X$ is not influenced by $U$ is standard to allow for classical IV regression and straightforward IV generation (Yuan et al., 2022; Li and Yao, 2024; Cheng et al., 2023; Chou et al., 2024). Since our feature decomposition produces both a confounder and an instrument, we add an additional constraint based on the following Lemma to ensure that $\widetilde{Z}$ is valid even when $X$ may be influenced by $U$.
Lemma 1. If $Z\sim\mathcal{N}(0,\sigma^{2})$, and suppose that the random variables $(Z,e_{Y},X,T)$ satisfy $\mbox{Cov}(Z,e_{Y}-\mathbb{E}[e_{Y}|X,T])=0$. Then it follows that $\mbox{Cov}(Z,e_{Y})=0$.
Proof. Notice that as $\mathbb{E}[Z]=0$, we have
$0$ $=\mbox{Cov}(Z,e_{Y}-\mathbb{E}[e_{Y}|X,T])$
$=\mathbb{E}[Z(e_{Y}-\mathbb{E}[e_{Y}|X,T])]-\mathbb{E}[Z]\,\mathbb{E}[(e_{Y}-\mathbb{E}[e_{Y}|X,T])]$
$=E[Z(e_{Y}-\mathbb{E}[e_{Y}|X,T])]$
$=\mathbb{E}[Z\cdot e_{Y}]-\mathbb{E}[Z]\mathbb{E}[e_{Y}|X,T]=\mbox{Cov}(Z,e_{Y}).\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\square$
This additional covariance term can be evaluated by fitting a regression model for $\hat{Y}=\mathbb{E}[Y|X,T]$ and then computing its residuals as $Y-\mathbb{E}[Y|X,T]$. Notice that $Y-\mathbb{E}[Y|X,T]=Y-\varphi(X,T)-(E[Y|X,T]-\varphi(X,T))=e_{Y}-\mathbb{E}[e_{Y}|X,T]$. Regardless of whether any existing instruments are normally distributed, if we can construct $g(X)=\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2})$ to have no covariance with the residuals $e_{Y}-\mathbb{E}[e_{Y}|X,T]$, i.e. with $\tilde{\varepsilon}_{Y}:=Y-\mathbb{E}[Y|X,T]$, then $\widetilde{Z}$ has zero covariance with $e_{Y}$ by Lemma 1. To sum up, $f$ and $g$ need to satisfy the following moment restrictions:
Moment Condition 1: $\mbox{Cov}(g(X),\tilde{\varepsilon}_{Y})=0$.
Moment Condition 2: $\mbox{Cov}(g(X),f(X))=0$.
Moment Condition 3: $\mbox{Cov}(f(X),Y)>0$.
Moment Condition 4: $\mbox{Cov}(T,g(X))>0$.
4 Related Work
Prior work on learning IVs has largely focused on the problem of IV selection from a set of observed candidates. For instance, ModeIV chooses instruments by looking at clusters of treatment effects based on weighting the observed variables used as instruments (Hartford et al., 2021). DIV-VAE uses a variational autoencoder (VAE) approach to disentangle an instrumental variable under the assumption that a surrogate instrument exists in the data (Cheng et al., 2024). IV-Tetrad (Silva and Shimizu, 2017) builds strong instruments requiring at least two observed IV candidates. Several methods for refining IV candidates and estimating causal effects from these candidates can fall under this category as well, including the sisVIVE (Kang et al., 2016), TEDVAE (Zhang et al., 2021) and Ivy models (Kuang et al., 2020) among various others (Davies et al., 2015; Burgess et al., 2016).
Methods to learn IVs were also proposed in prior work, including GIV (Wu et al., 2023), AutoIV (Yuan et al., 2022), VIV (Li and Yao, 2024), DVAE-CIV (Cheng et al., 2023), and GDIV (Chou et al., 2024). GIV generates a categorical IV with unsupervised expectation-maximization that groups data according to underlying distributional differences assumed to arise from the aggregation of data from multiple sources. This approach uses environment as an IV in data coming multiple sources (Schweisthal et al., 2024). The other methods learn variational distributions. The AutoIV method uses a mutual information (MI) based loss to generate an abstract IV from observed data by learning variational distributions (Yuan et al., 2022). VIV, DVAE-CIV, and GDIV use VAEs to learn independent latent variables that serve as $Z,U,C$ from the observed data $Y,T,X$, sometimes including an additional adjustment variable $A$ derived from $Y,X$. VAEs have shown great success in probabilistic modeling in general but lack theory to guarantee learning the true causal model and satisfaction of IV conditions.
5 The ZNet Model
5.1 ZNet Architecture
The ZNet architecture encodes an inductive prior consistent with the SCM of the IV setup in (1). The ZNet model comprises two encoders, $f$ and $g$, which learn representations of the confounders $\widetilde{X}$ and the instrument $\widetilde{Z}$, respectively, along with three feedforward neural networks, $\Phi,\varphi$ and $\pi$, used to estimate $\mathbb{E}[Y|X,T],\mathbb{E}[Y|\widetilde{X},T]$ and $\mathbb{E}[T|\widetilde{Z}$]. An overview of all components of the ZNet model is shown in Fig. 3. After training, ZNet produces embeddings $\{\widetilde{X},\widetilde{Z}\}$ from the observed features $X$, which we then use to transform the observational dataset $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i}$ into a new dataset $\widetilde{\mathcal{D}}=\{(\widetilde{X}_{i},\widetilde{Z}_{i},T_{i},Y_{i})\}_{i}$ with the learned instruments $\{\widetilde{Z}_{i}\}_{i}$. Importantly, as we show in the experiments section, ZNet is compatible with any downstream IV-based causal estimator.
Figure 3: Overview of the components of the ZNet Architecture.
5.2 ZNet Loss Function
[…]
$\min_{f,g,\varphi,\pi}\mathcal{L}_{Y}(\varphi(f))+\mathcal{L}_{T}(\pi(g))\,\,\text{s.t.}\,\begin{cases}\operatorname{Cov}(g(X),\tilde{\varepsilon}_{Y})=0,\\ \operatorname{Cov}(f(X),g(X))=0.\end{cases}$
Here, $\mathcal{L}_{Y}(\varphi(f))$ and $\mathcal{L}_{T}(\pi(g))$ are supervised losses for the models $\varphi(f(X),T)$ and $\pi(g(X))$ against the labels $Y$ and $T$ (e.g., mean squared error and cross entropy losses). Note that moment conditions 3 and 4 in Section 3 are implicit in these losses. We train ZNet through a Lagrangian relaxation of the constrained optimization problem above as follows:
$\min_{f,g,\varphi,\pi}\mathcal{L}_{Y}(\varphi(f))+\mathcal{L}_{T}(\pi(g))$ $+\lambda_{0}\operatorname{Cov}(g(X),\tilde{\varepsilon}_{Y})$
$+\lambda_{1}\operatorname{Cov}(f(X),g(X)),$
where $\lambda_{0}$ and $\lambda_{1}$ are tunable hyperparameters. Our training algorithm approximates the covariance terms above using a weighted combination of Pearson correlation coefficients or mutual information, as described in the next section.
5.3 ZNet Implementation and Training Details
The networks $\Phi,\varphi,\pi,f,g$ are implemented as feedforward networks, where the activation function can be chosen between ReLU or linear. Supervised learning losses for $\Phi,\varphi$ and $\pi$ are implemented as mean squared error (MSE), MSE, and binary cross entropy (BCE), respectively.
[…]
$\mathcal{L}_{\mbox{ZNet}}$ $=\alpha_{1}\operatorname{MSE}(\varphi(f(X),T),Y)$
$+\alpha_{2}\operatorname{BCE}(\pi(g(X)),T)$
$+\alpha_{3}(1-\widehat{\operatorname{Cov}}(\varphi(f(X),T),Y)^{2})$
$+\alpha_{4}(1-\widehat{\operatorname{Cov}}(\pi(g(X)),T)^{2})$
$+\alpha_{5}\widehat{\operatorname{Cov}}(g(X),\tilde{\varepsilon}_{Y})^{2}+\alpha_{6}\widehat{\operatorname{Cov}}(\pi(g(X)),T)^{2}$
$+\alpha_{7}\operatorname{KL}(g(X),\mathcal{N})+\alpha_{8}\operatorname{KL}(f(X),\mathcal{N})$
$+\alpha_{9}\widehat{\operatorname{Cov}}(g(X))^{2}+\alpha_{10}\widehat{\operatorname{Cov}}(f(X))^{2}$
$+\alpha_{11}\operatorname{MSE}(\Phi(f(X)),Y).$
Training steps. ZNet training occurs in three stages. First, the network $\Phi$ is separately trained to estimate $\tilde{\varepsilon}_{Y}$, i.e. only $\alpha_{11}$ is non-zero. Weights of $\Phi$ are then frozen. Next the full network is pretrained by fitting $\varphi,\pi$ in a supervised setting without the moment constraints ($\alpha_{5}=\alpha_{6}=\alpha_{7}=\alpha_{8}=\alpha_{9}=\alpha_{10}=\alpha_{11}=0$). This serves to encourage starting representations of $\tilde{X}$ and $\tilde{Z}$ to be highly correlated with $Y$ and $T$, respectively. Finally, the full ZNet model is finetuned end-to-end across all the loss terms above (with $\alpha_{11}=0$).
Hyperparameter tuning. All hyperparameters, including all loss term weights and whether the moment constraints are approximated via Pearson correlations or MI, were tuned using Bayesian optimization implemented in Botorch (Balandat et al., 2020). We perform optimization corresponding to the two stages of IV regression. For each IV generation, we maximize the instrument’s relevance F-Statistic and minimize the correlation between learned $\tilde{X}$ and using Botorch’s native adaptation of the Noisy Expected Improvement acquisition function for multi-objective optimization. We then choose the parameter set from the Pareto front with the highest F-Statistic. We tune the treatment effect estimators to simultaneously minimize the MSE of the model’s ATE against a nearest-neighbors (NN) ATE and the MSE of estimated factual $Y$, again with the Noisy Expected Improvement acquisition function. The parameter set is selected from the Pareto front by the least MSE of NN.
5.4 Causal Effect Estimation
Recall that ZNet creates a dataset with learned confounders and instruments, and is therefore compatible with a wide range of downstream IV-based estimators of causal effects. We wrap our ZNet model around three downstream estimators of treatment effects to demonstrate the utility of ZNet in causal inference: TSLS, DFIV and DeepIV. Each method takes as input the true treatment term $T$ and our learned representations $\tilde{X}$ and $\tilde{Z}$. TSLS is the classical IV estimator. It assumes linear structural equations and independence of $U$ and $X$, thus we omit when not applicable (Imbens and Rubin, 2015; Verbeek, 2004). DeepIV (Hartford et al., 2017) generalizes TSLS by allowing the model at each stage to be parameterized by a neural network and $X\not\perp U$. The DFIV estimator allows basis functions at each stage to be parametrized by neural networks (Xu et al., 2020).
6 Experiments
Datasets. We conduct a comprehensive and extensive evaluation experiments of ZNet using the semi-synthetic IHDP dataset, a widely used benchmark for causal inference (Hill, 2011). IHDP is based on an experiment that studied the effect of home visits during infancy on cognitive test scores of premature infants, with 985 individuals and 25 covariates. We design multiple experimnetla setups using IHDP that correspond to different ways in which IVs may explicitly or implicitly exists. Each setup is define by three sets of features, $X^{\to T}$, $X^{\to Y}$, and $X^{\leftarrow U}$, where we each $X^{\to I}$ is the subset of covariates $X$ which have causal influence on $I$ in the arrow’s direction. We create four experimental setups based on their inclusion of an instrument:
[…]
Figure 4: Regression plots for learned and true instruments. Learned instruments are correlated to true ones in the setup with true instrument candidate. Here, $X_{13},X_{14},X_{15}$ are true IVs ($x$-axis) plotted against the corresponding dimensions of the learned $\widetilde{Z}$ ($y$-axis).
Directed acyclic graphs for all settings above are included in the Appendix. For each setting, we consider scenarios where $X^{\leftarrow U}\neq\varnothing$ and $X^{\leftarrow U}=\varnothing$. We also consider data where $U=\varnothing$ (i.e. no unobserved confounding). After fixing a covariate set, we choose functions $\phi,\psi,e_{Y},e_{T}$ and generate the variables $Y,T$ similar to (Wu et al., 2023) by writing
$Y$ $=\phi(X_{Y},T)+e_{Y}(U)+\varepsilon_{Y}\mbox{ for }\varepsilon_{Y}\sim\mathcal{N}(0,.1),$
$T$ $\sim\mbox{Bernoulli}(P)\mbox{ for }P=\psi(X_{T})+e_{T}(U)+\varepsilon_{T},$
for $\varepsilon_{T}\sim\mathcal{N}(0,.1)$. We focus on binary treatments, though ZNet is easily adapted for continuous settings. We consider linear and non-linear version of $\phi,\psi$ for each dataset. For each scenario, we randomly generate 10 datasets with varying functions $\phi,e_{Y},\varepsilon_{Y},\psi,e_{T},\varepsilon_{T}$. Across dataset and confounding settings, this means we evaluate ZNet on a comprehensive collection of 180 datasets. Data are always split into 60% for training, 20% for tuning, and 20% for testing. All experiments report results on the test proportion.
Metrics and baselines. All model performances were evaluated using bootstrapping (N=50). ATE errors are evaluated in terms of absolute proportions: $|\mbox{Estimate}-\mbox{True}|/|\mbox{True}|$. Error on CATE estimation is presented as average square error in estimated individual level effect (PEHE). In addition to prior methods discussed in Section 4, we compare ZNet-learned instruments to using the true instrument TrueIV, if it exists, and to TARNet (Shalit et al., 2017), a CATE estimation model used in the absence of unobserved confounding.
Table 1: Evaluating the average instrumental properties of ZNet learned representations
Metric ZNet TrueIV
(Train / Val / Test) (Train / Val / Test)
F-Stat($\tilde{Z}$,$T$) 140.25 / 14.82 / 8.70 57.36 / 19.35 / 8.18
Corr($\tilde{Z}$$\tilde{X}$) 0.15 / 0.16 / 0.19 0.09 / 0.10 / 0.16
Corr($\tilde{Z}$,$U$) 0.13 / 0.13 / 0.15 0.03 / 0.06 / 0.06
Corr($\tilde{Z}$, $\tilde{\varepsilon}_{Y}$) 0.02 / 0.06 / 0.09 0.04 / 0.05 / 0.08
6.1 ZNet Instrument Representations
As shown in Fig. 4, ZNet successfully recovers existing instruments. In the Linear Mixed Candidate setting, there are three variables $X_{13},X_{14},X_{15}\in X^{\to T}$ which are explicit instruments. In one selected sample dataset, ZNet generates a 10-dimensional variable $Z$ which is correlated with and linearly predicts each of $X_{13},X_{14},X_{15}$. ZNet also recovers latent instruments. Consider an example Linear Categorical Instrument dataset. The true instrument groups the observed data into 5 clusters. ZNet approximately recovers these clusters by regression, using t-SNE dimensionality reduction, and K-Means with cluster relabeling in Fig. 5.
Table 2: Mean absolute ATE proportional error (SE) across all experimental settings with the DeepIV and DF-IV estimators.
[…]
TARNet 180 0.442 (0.03)
DeepIV TrueIV 120 0.406 (0.03)
ZNet 180 0.412 (0.03)
AutoIV 0.440 (0.03)
VIV 0.438 (0.04)
GIV 0.429 (0.03)
DF IV TrueIV 120 0.604 (0.09)
ZNet 180 0.544 (0.04)
AutoIV 0.705 (0.06)
VIV 0.652 (0.11)
GIV 0.798 (0.07)
Figure 5: Learned instrument representations are correlated to existing instruments in latent categorical instrument dataset. Left most plot are regression predictions of the true categorical IV using learned $\widetilde{Z}$. Middle plot is a t-SNE visualization of $\widetilde{Z}$ colored by the true IV. Right most is a normalized confusion matrix of K-means clusters of $\widetilde{Z}$ compared to the true instruments.
Independent of the existence of an instrument in the observed data, the ZNet model generates a representation that is correlated with $T$, independent of the confounder representation $\widetilde{X}$, independent of the error in predicting $Y$, and unconfounded by $U$. We evaluate the suitability of these instrument representations empirically. We observe F-Statistics that indicate the predictive power of generated representations $\widetilde{Z}$ for $T$ and low correlation across prohibited relationships between $\widetilde{Z}$ and confounders. ZNet IV condition satisfaction is strong on average over all 180 datasets compared to the average over existing dataset IVs (Table 1). We additionally compare satisfaction by dataset scenarios in Appendix Tables 11, 12.
6.2 Causal Effect Estimation with ZNet
We used the ZNet learned instrument representations to recover the causal effects using a second stage regression with DeepIV or DFIV. We assess performance across the all experimental settings. We highlight performance across three comparisons: (1) all 180 datasets, (2) the 80 datasets with confounding influencing $X$, i.e. $U\to X$ datasets, and (3) the 40 confounded datasets with no instrument candidate.
The first comparison shows that the performance of ZNet in estimating ATE outperforms baselines across the different experimental settings (Table 2). Notably, ZNet is the only method that outperforms TARNet, which ignores unobserved confounding altogether. Additionally, ZNet performs best in the setting where the unobserved confounding influences the observed data, i.e. $U\to X$ datasets (Table 3). This is because our loss construction guarantees this owing to the inclusion of the first moment condition (Lemma 1).
Table 3: Mean absolute ATE proportional error (SE) across $U\to X$ datasets with the DeepIV and DF-IV estimators.
[…]
TARNet 80 0.666 (0.06)
DeepIV TrueIV 60 0.534 (0.05)
ZNet 80 0.620 (0.06)
AutoIV 0.677 (0.06)
VIV 0.702 (0.07)
GIV 0.672 (0.06)
DF IV TrueIV 60 0.736 (0.17)
ZNet 80 0.729 (0.08)
AutoIV 0.870 (0.10)
VIV 0.906 (0.15)
GIV 0.860 (0.09)
The performance comparison in Table 4 focuses on the most common setting where explicit instrument candidates are not available. ZNet once again outperforms existing IV generation methods and TARNet. Performance on causal effect estimation was worse using DF IV than DeepIV consistently regardless of data generation method. We also include additional results stratified by the data generation settings which show where IV estimation is easier or more challenging and results of analyses with TSLS in the appendix (See Appendix Tables 7, 8). We report results on the PEHE in CATE estimation, which similarly varies by data generation settings, in the appendix as well (Appendix Tables 9, 10).
Table 4: Mean absolute ATE proportional error (SE) across confounded settings with no true instrument candidate.
[…]
TARNet 40 0.619 (0.10)
DeepIV ZNet 40 0.566 (0.09)
AutoIV 0.667 (0.10)
VIV 0.654 (0.11)
GIV 0.704 (0.11)
DF IV ZNet 40 0.773 (0.14)
AutoIV 0.837 (0.14)
VIV 0.986 (0.28)
GIV 1.077 (0.17)
6.3 Application of ZNet to Unstructured Data
ZNet in the unstructured data setting is particularly powerful because high dimensional data can contain additional information recoverable as an instrument more readily. We believe this is especially impactful in the setting of unstructured health data. We demonstrate the ability of ZNet to recover instruments in such settings using a semi-synthetic electrocardiogram (ECG) dataset. Data from 35,463 ECGs from a public Physionet dataset serve as $X$, age and sex serve as hidden confounders $U$, and variables $T$ and $Y$ are sampled, analogous to the process for IHDP data, in this case as a random linear combination of eight interpretable, derived features of the ECG, i.e. heart rate, QT interval, QRS amplitude, etc. (Zheng et al., 2022, 2000; Goldberger et al., 2000). As with most unstructured health data, no true instrument exists. Despite that, ZNet still recovers an instrument representation and more accurately predicts treatment effects than ordinary least squares regression using the tabular features (Table 5). Here, we use ZNet with Deep IV since it is consistently more accurate than DF-IV.
Table 5: Mean absolute ATE proportional error (SE) across bootstraps in the ECG dataset.
[…]
DeepIV ZNet 0.126 (0.000647)
OLS 0.665 (0.00263)
These ECG data come from two hospitals in Zhejiang, China. In true observational data, both treatments received and images taken can vary by care setting due to purchased equipment or physician propensity despite being random with respect to many confounders. Moreover, the outcomes associated to treatments are typically uncorrelated with hospital setting directly. While hospital setting is not available in the ECG data and its instrumental value hidden to an analyst, with ZNet, we extract a representation that satisfies instrument properties potentially leveraging this latent information as illustrated pictorially in Fig. 6. We evaluate the instrument selected by ZNet by its satisfaction of IV properties as correlations in Table 6. Values suggest a relevant instrument that through weak correlations in prohibited relations also satisfies exclusion restriction and unconfoundedness.
Table 6: Evaluating the IV conditions of the ECG based instrument.
Metric ZNet ECG IV
(Train / Val / Test)
F-Stat(Z~,T)($\tilde${Z},T) 261.26 / 59.69 / 22.57
[…]
$\mbox{Cor}(\tilde{Z},U_{1})$ 0.25 / 0.25 / 0.244
$\mbox{Cor}(\tilde{Z},U_{2})$ -0.15 / -0.16 / -0.15
$\mbox{Cor}(\tilde{Z},\tilde{\varepsilon}_{Y})$ 0.0022 / 0.0048 / 0.015
7 Discussion
Figure 6: Visualization of ZNet with unstructured health data.
ZNet enables IV regression without domain knowledge of pre-existing IVs. This automation of IV generation from observed data is exciting, as it can enable the widespread use of IV regression for causal effect estimation. To this end, we contribute a comprehensive evaluation of IV generation for causal inference, which demonstrates broad utility. Regardless of the existence of a candidate or a latent instrument or of unobserved confounding, ZNet can exceed the performance of TARNet and of probabilistic IV generation methods. This performance is enabled by the recovery of valid instrument representations. In the case of existing instruments among the observed data, we see that recovered instruments are highly correlated with these variables. Without candidates, ZNet representations are still relevant and show minimal correlation with hidden confounders and directly with the outcome.
The data generation process of any real world dataset is untestable. We can never confirm a lack of unobserved confounding, even when evaluating a candidate instrument. ZNet’s strong performance across settings suggests that it can serve as a “plug-in" causal inference estimator for causal inference in general observational settings, regardless of whether the untestable assumption of unconfoundedness is satisfied. Our evaluations show ZNet to be particularly well suited when unobserved confounders influence the observed data and where instrument candidates do not exist. These are the most challenging and general settings.
ZNet is especially relevant in the context of high dimensional unstructured data. Unstructured data may contain latent or abstract instruments more frequently, as high-dimensional feature spaces often contain rich information that our approach could learn to extract as instruments. High dimensional tabular and image data are becoming ubiquitous in the predictive models used across industries. Interest grows in embedding causal relationships into these models so that insights from their features are actionable. ZNet automatically mitigates unobserved confounding broadening the ability to pursue causal machine learning in predictive settings. We illustrate ZNet’s treatment effect estimation capability with ECGs, but applications for ZNet extend far beyond this modality and beyond healthcare.
Solutions to the ZNet loss minimization problem always give a representation that serves empirically as an instrument since empirical IV constraints are explicitly embedded in the loss function. This instrument can then be used in any downstream IV regression where satisfying the standard IV criteria (or, equivalently, ZNet moment constraints) implies the validity of subsequent causal inference. The strength and validity of ZNet generated instruments vary in finite settings. Variance in finite samples affects our ability to predict treatment effects. However, IV estimation is also limited in general by a lack of theoretical guarantees of identifiably without conditions on the SCM. This theoretically limits not only our approach but IV estimation in general. Still, strong empirical results and opportunities in high dimensional data suggest the value in the increased use of these methods beyond linear settings.
Code
All code for running ZNet is available at https://github.com/AlaaLab/ZNet.git.
Appendix A Additional Figures
Table 7: ATE Results on Synthetic Linear Datasets
Dataset Diff Means TARNet TSLS IV Method DeepIV DF IV
Linear Disjoint 0.318 0.140 0.210 TrueIV 0.112 0.175
ZNet 0.110 0.229
AutoIV 0.090 0.329
VIV 0.088 0.184
GIV 0.113 0.537
Linear Disjoint (no $U\to X$) 0.267 0.207 0.327 TrueIV 0.148 0.280
ZNet 0.222 0.469
AutoIV 0.199 0.775
VIV 0.190 0.269
GIV 0.168 0.521
Linear Latent Categorical 0.160 0.157 1.118 TrueIV 0.101 0.171
ZNet 0.090 0.080
AutoIV 0.134 0.626
VIV 0.092 0.353
GIV 0.095 0.492
Linear Latent Categorical (no $U\to X$) 0.390 0.372 1.088 TrueIV 0.389 0.384
ZNet 0.324 0.453
AutoIV 0.343 0.723
VIV 0.309 0.411
GIV 0.298 0.535
Linear Mixed 0.790 0.570 0.448 TrueIV 0.660 0.592
ZNet 0.537 0.844
AutoIV 0.690 0.886
VIV 0.716 0.859
GIV 0.685 0.750
Linear Mixed (no $U\to X$) 0.390 0.298 0.326 TrueIV 0.314 0.446
ZNet 0.314 0.535
AutoIV 0.344 0.463
VIV 0.375 0.411
GIV 0.370 0.324
Linear No Candidate 1.749 1.375 TrueIV
ZNet 1.306 1.724
AutoIV 1.418 1.630
VIV 1.460 1.429
GIV 1.491 1.546
Linear No Candidate (no $U\to X$) 0.291 0.365 TrueIV
ZNet 0.361 0.713
AutoIV 0.405 0.831
VIV 0.367 0.499
GIV 0.435 1.193
Linear No Candidate (no $U$) 0.060 0.092 TrueIV
ZNet 0.037 0.245
AutoIV 0.064 0.823
VIV 0.043 0.190
GIV 0.052 0.556
Table 8: ATE Results on Synthetic Non-Linear Datasets
Dataset Diff Means TARNet TSLS IV Method DeepIV DF IV
Non-linear Disjoint 0.845 0.571 0.424 TrueIV 0.586 0.602
ZNet 0.553 0.706
AutoIV 0.656 0.869
VIV 0.633 0.558
GIV 0.551 0.854
Non-linear Disjoint (no $U\to X$) 0.172 0.255 0.349 TrueIV 0.202 0.661
ZNet 0.274 0.354
AutoIV 0.283 0.606
VIV 0.218 0.339
GIV 0.215 1.951
Non-linear Latent Categorical 1.441 0.784 12.998 TrueIV 0.697 1.758
ZNet 0.689 0.784
AutoIV 0.581 1.080
VIV 0.798 0.678
GIV 0.647 0.727
Non-linear Latent Categorical (no $U\to X$) 0.431 0.405 1.569 TrueIV 0.375 0.499
ZNet 0.350 0.283
AutoIV 0.376 0.319
VIV 0.331 0.253
GIV 0.299 0.810
Non-linear Mixed 1.501 1.190 1.570 TrueIV 1.051 1.118
ZNet 1.200 1.005
AutoIV 1.156 0.971
VIV 1.179 1.397
GIV 1.098 0.938
Non-linear Mixed (no $U\to X$) 0.205 0.342 0.535 TrueIV 0.232 0.564
ZNet 0.379 0.327
AutoIV 0.250 0.351
VIV 0.245 1.786
GIV 0.254 0.635
Non-linear No Candidate 0.658 0.542 TrueIV
ZNet 0.478 0.457
AutoIV 0.690 0.567
VIV 0.654 1.788
GIV 0.698 1.040
Non-linear No Candidate (no $U\to X$) 0.216 0.193 TrueIV
ZNet 0.121 0.200
AutoIV 0.155 0.321
VIV 0.133 0.229
GIV 0.190 0.528
Non-linear No Candidate (no $U$) 0.109 0.105 TrueIV
ZNet 0.064 0.387
AutoIV 0.079 0.513
VIV 0.046 0.110
GIV 0.068 0.433
Table 9: PEHE on Linear Synthetic Dataset.
Dataset TARNet IV Method DeepIV DF IV
Linear Disjoint 0.142 TrueIV 0.165 0.187
ZNet 0.189 0.240
AutoIV 0.150 0.323
VIV 0.153 0.209
GIV 0.188 0.445
Linear Disjoint (no $U\to X$) 0.176 TrueIV 0.202 0.272
ZNet 0.203 0.431
AutoIV 0.224 0.541
VIV 0.215 0.273
GIV 0.237 0.428
Linear Latent Categorical 0.213 TrueIV 0.264 0.259
ZNet 0.239 0.212
AutoIV 0.283 0.692
VIV 0.263 0.470
GIV 0.230 0.627
Linear Latent Categorical (no $U\to X$) 0.421 TrueIV 0.524 0.409
ZNet 0.477 0.502
AutoIV 0.508 0.676
VIV 0.495 0.451
GIV 0.471 0.501
Linear Mixed 0.681 TrueIV 0.805 0.800
ZNet 0.692 1.167
AutoIV 0.828 0.943
VIV 0.836 0.969
GIV 0.824 0.855
Linear Mixed (no $U\to X$) 0.385 TrueIV 0.527 0.672
ZNet 0.593 0.817
AutoIV 0.566 0.658
VIV 0.599 0.536
GIV 0.540 0.493
Linear No Candidate 0.726 TrueIV
ZNet 0.741 0.935
AutoIV 0.799 0.793
VIV 0.825 0.777
GIV 0.799 0.801
Linear No Candidate (no $U\to X$) 0.428 TrueIV
ZNet 0.464 0.912
AutoIV 0.563 0.884
VIV 0.530 0.587
GIV 0.568 1.321
Linear No Candidate (no $U$) 0.178 TrueIV
ZNet 0.229 0.570
AutoIV 0.342 1.553
VIV 0.224 0.452
GIV 0.313 0.932
Table 10: PEHE on Non-linear Synthetic Dataset.
Dataset TARNet IV Method DeepIV DF IV
Non-linear Disjoint 0.788 TrueIV 0.762 0.942
ZNet 0.733 0.855
AutoIV 0.805 1.028
VIV 0.818 0.700
GIV 0.753 0.930
Non-linear Disjoint (no $U\to X$) 0.548 TrueIV 0.509 0.834
ZNet 0.528 1.253
AutoIV 0.556 0.846
VIV 0.523 0.567
GIV 0.509 2.245
Non-linear Latent Categorical 0.529 TrueIV 0.561 1.121
ZNet 0.515 0.548
AutoIV 0.519 0.954
VIV 0.626 0.552
GIV 0.524 0.520
Non-linear Latent Categorical (no $U\to X$) 0.430 TrueIV 0.493 0.472
ZNet 0.436 0.304
AutoIV 0.527 0.397
VIV 0.455 0.303
GIV 0.446 0.709
Non-linear Mixed 0.965 TrueIV 0.818 0.786
ZNet 0.988 0.866
AutoIV 0.956 0.777
VIV 1.005 0.882
GIV 0.887 0.704
Non-linear Mixed (no $U\to X$) 0.952 TrueIV 0.842 0.963
ZNet 0.854 0.859
AutoIV 0.841 0.846
VIV 0.858 2.192
GIV 0.820 1.245
Non-linear No Candidate 0.648 TrueIV
ZNet 0.624 0.615
AutoIV 0.730 0.622
VIV 0.724 1.417
GIV 0.709 0.822
Non-linear No Candidate (no $U\to X$) 0.833 TrueIV
ZNet 0.818 0.802
AutoIV 0.820 0.801
VIV 0.805 0.775
GIV 0.793 1.249
Non-linear No Candidate (no $U$) 1.165 TrueIV
ZNet 1.081 1.393
AutoIV 1.158 1.415
VIV 1.181 1.070
GIV 1.144 1.146
Table 11: Instrument Strength and Validity Linear Synthetic Dataset.
Dataset IV Method F-Stat(Z,T) (Relevance) (Train/Val/Test) Corr(Z,C) (Independence) (Train/Val/Test) Corr(Z,Y-Yhat) (Exogeneity) (Train/Val/Test) Corr(Z,U) (Independence) (Train/Val/Test)
Linear Disjoint TrueIV 9.792 / 3.215 / 3.657 0.027 / 0.051 / 0.124 0.032 / 0.069 / 0.095 0.030 / 0.062 / 0.045
ZNet 28.439 / 2.444 / 3.085 0.128 / 0.141 / 0.180 0.033 / 0.068 / 0.090 0.158 / 0.166 / 0.170
AutoIV 5.275 / 1.880 / 2.731 0.208 / 0.221 / 0.243 0.000 / 0.000 / 0.000 0.104 / 0.109 / 0.098
VIV 17.719 / 5.397 / 4.568 0.032 / 0.057 / 0.112 0.032 / 0.049 / 0.101 0.036 / 0.050 / 0.075
GIV 8.737 / 2.217 / 3.190 0.139 / 0.143 / 0.173 0.032 / 0.039 / 0.067 0.080 / 0.084 / 0.111
Linear Disjoint (no $U\to X$) TrueIV 17.167 / 7.477 / 6.471 0.027 / 0.051 / 0.124 0.051 / 0.056 / 0.116 0.035 / 0.049 / 0.095
ZNet 45.214 / 8.188 / 3.851 0.080 / 0.111 / 0.149 0.025 / 0.052 / 0.095 0.029 / 0.056 / 0.093
AutoIV 9.318 / 2.783 / 2.159 0.212 / 0.222 / 0.255 0.000 / 0.000 / 0.000 0.030 / 0.062 / 0.091
VIV 16.739 / 3.926 / 3.641 0.030 / 0.058 / 0.119 0.025 / 0.065 / 0.109 0.036 / 0.053 / 0.080
GIV 0.913 / 1.251 / 2.414 0.148 / 0.149 / 0.190 0.024 / 0.045 / 0.072 0.039 / 0.062 / 0.087
Linear Latent Categorical TrueIV 165.204 / 53.555 / 19.690 0.202 / 0.206 / 0.230 0.011 / 0.014 / 0.028 0.035 / 0.048 / 0.076
ZNet 179.695 / 25.554 / 9.951 0.134 / 0.134 / 0.165 0.014 / 0.039 / 0.071 0.074 / 0.081 / 0.098
AutoIV 68.529 / 18.436 / 6.537 0.258 / 0.247 / 0.263 0.000 / 0.000 / 0.000 0.144 / 0.154 / 0.156
VIV 17.504 / 5.266 / 5.422 0.035 / 0.064 / 0.118 0.030 / 0.039 / 0.097 0.036 / 0.056 / 0.082
GIV 97.670 / 25.682 / 12.462 0.140 / 0.157 / 0.178 0.020 / 0.024 / 0.065 0.036 / 0.067 / 0.081
Linear Latent Categorical (no $U\to X$) TrueIV 199.602 / 65.321 / 24.395 0.204 / 0.212 / 0.235 0.006 / 0.016 / 0.031 0.036 / 0.049 / 0.065
ZNet 300.974 / 25.263 / 11.940 0.111 / 0.117 / 0.138 0.010 / 0.053 / 0.085 0.043 / 0.058 / 0.073
AutoIV 91.941 / 28.877 / 8.274 0.249 / 0.263 / 0.266 0.000 / 0.000 / 0.000 0.031 / 0.054 / 0.074
VIV 16.884 / 4.987 / 4.814 0.034 / 0.058 / 0.115 0.024 / 0.042 / 0.099 0.031 / 0.064 / 0.081
GIV 35.204 / 24.341 / 9.507 0.146 / 0.146 / 0.191 0.017 / 0.038 / 0.056 0.029 / 0.054 / 0.082
Linear Mixed TrueIV 8.070 / 2.816 / 3.181 0.027 / 0.051 / 0.124 0.081 / 0.099 / 0.113 0.030 / 0.062 / 0.045
ZNet 221.401 / 25.663 / 20.673 0.209 / 0.208 / 0.238 0.013 / 0.053 / 0.084 0.284 / 0.240 / 0.303
AutoIV 103.603 / 22.816 / 23.096 0.209 / 0.211 / 0.235 0.000 / 0.000 / 0.000 0.253 / 0.219 / 0.289
VIV 18.179 / 5.523 / 4.300 0.033 / 0.053 / 0.111 0.030 / 0.044 / 0.112 0.050 / 0.072 / 0.091
GIV 43.053 / 8.902 / 8.220 0.147 / 0.150 / 0.195 0.026 / 0.041 / 0.080 0.124 / 0.117 / 0.157
Linear Mixed (no $U\to X$) TrueIV 12.066 / 5.101 / 3.489 0.027 / 0.051 / 0.124 0.063 / 0.081 / 0.097 0.027 / 0.060 / 0.079
ZNet 133.538 / 16.742 / 8.084 0.150 / 0.155 / 0.186 0.025 / 0.077 / 0.114 0.045 / 0.055 / 0.094
AutoIV 104.522 / 31.072 / 17.727 0.237 / 0.241 / 0.261 0.000 / 0.000 / 0.000 0.029 / 0.049 / 0.094
VIV 16.356 / 7.037 / 3.971 0.033 / 0.060 / 0.117 0.028 / 0.053 / 0.101 0.042 / 0.055 / 0.086
GIV 12.316 / 2.874 / 4.237 0.150 / 0.159 / 0.194 0.027 / 0.026 / 0.063 0.029 / 0.042 / 0.095
Linear No Candidate TrueIV
ZNet 141.454 / 17.320 / 8.324 0.150 / 0.168 / 0.199 0.019 / 0.070 / 0.079 0.327 / 0.310 / 0.295
AutoIV 88.696 / 23.879 / 11.640 0.245 / 0.240 / 0.272 0.000 / 0.000 / 0.000 0.404 / 0.399 / 0.373
VIV 18.248 / 4.797 / 4.426 0.033 / 0.059 / 0.115 0.039 / 0.062 / 0.098 0.057 / 0.074 / 0.098
GIV 52.157 / 15.182 / 5.950 0.145 / 0.141 / 0.186 0.029 / 0.039 / 0.075 0.124 / 0.144 / 0.134
Linear No Candidate (no $U\to X$) TrueIV
ZNet 71.331 / 2.770 / 3.064 0.166 / 0.181 / 0.213 0.024 / 0.069 / 0.095 0.044 / 0.060 / 0.079
AutoIV 16.775 / 2.015 / 2.158 0.236 / 0.242 / 0.254 0.000 / 0.000 / 0.000 0.041 / 0.057 / 0.085
VIV 18.499 / 7.769 / 4.560 0.032 / 0.065 / 0.115 0.037 / 0.052 / 0.099 0.037 / 0.068 / 0.072
GIV 0.850 / 0.970 / 1.879 0.154 / 0.166 / 0.207 0.020 / 0.033 / 0.066 0.028 / 0.058 / 0.084
Linear No Candidate (no $U$) TrueIV
ZNet 115.223 / 2.995 / 2.728 0.147 / 0.156 / 0.190 0.031 / 0.062 / 0.082 – / – / –
AutoIV 12.631 / 2.437 / 2.154 0.216 / 0.232 / 0.250 0.000 / 0.000 / 0.000 – / – / –
VIV 21.595 / 5.473 / 4.768 0.031 / 0.054 / 0.116 0.037 / 0.047 / 0.113 – / – / –
GIV 1.574 / 0.649 / 1.787 0.161 / 0.164 / 0.200 0.022 / 0.018 / 0.062 – / – / –
Table 12: Instrument Strength and Validity on Non-linear Synthetic Dataset.
Dataset IV Method F-Stat(Z,T) (Relevance) (Train/Val/Test) Corr(Z,C) (Independence) (Train/Val/Test) Corr(Z,Y-Yhat) (Exogeneity) (Train/Val/Test) Corr(Z,U) (Independence) (Train/Val/Test)
Non-linear Disjoint TrueIV 10.436 / 4.195 / 2.690 0.027 / 0.051 / 0.124 0.079 / 0.078 / 0.108 0.030 / 0.062 / 0.045
ZNet 298.079 / 24.161 / 14.563 0.163 / 0.178 / 0.192 0.031 / 0.049 / 0.080 0.285 / 0.256 / 0.257
AutoIV 102.265 / 24.380 / 18.903 0.206 / 0.217 / 0.227 0.000 / 0.000 / 0.000 0.278 / 0.241 / 0.295
VIV 16.532 / 5.867 / 5.673 0.033 / 0.058 / 0.114 0.034 / 0.047 / 0.094 0.059 / 0.069 / 0.102
GIV 17.338 / 5.966 / 5.895 0.141 / 0.150 / 0.181 0.036 / 0.048 / 0.074 0.144 / 0.140 / 0.176
Non-linear Disjoint (no $U\to X$) TrueIV 19.393 / 6.986 / 5.568 0.027 / 0.051 / 0.124 0.028 / 0.055 / 0.115 0.030 / 0.044 / 0.094
ZNet 70.675 / 7.622 / 4.336 0.143 / 0.157 / 0.174 0.017 / 0.053 / 0.105 0.033 / 0.046 / 0.081
AutoIV 40.857 / 6.378 / 3.899 0.236 / 0.236 / 0.260 0.000 / 0.000 / 0.000 0.031 / 0.043 / 0.090
VIV 17.436 / 5.063 / 4.857 0.034 / 0.056 / 0.117 0.042 / 0.051 / 0.101 0.034 / 0.058 / 0.076
GIV 2.574 / 1.176 / 1.936 0.140 / 0.148 / 0.187 0.042 / 0.039 / 0.067 0.035 / 0.060 / 0.072
Non-linear Latent Categorical TrueIV 96.409 / 39.441 / 8.684 0.204 / 0.212 / 0.235 0.009 / 0.011 / 0.025 0.064 / 0.083 / 0.022
ZNet 120.953 / 20.441 / 7.507 0.173 / 0.172 / 0.188 0.030 / 0.069 / 0.099 0.174 / 0.178 / 0.179
AutoIV 58.730 / 12.420 / 9.422 0.213 / 0.221 / 0.236 0.000 / 0.000 / 0.000 0.251 / 0.231 / 0.250
VIV 15.337 / 3.719 / 4.762 0.034 / 0.061 / 0.113 0.034 / 0.043 / 0.091 0.037 / 0.063 / 0.083
GIV 57.630 / 20.981 / 10.098 0.131 / 0.142 / 0.169 0.025 / 0.035 / 0.068 0.118 / 0.103 / 0.149
Non-linear Latent Categorical (no $U\to X$) TrueIV 129.011 / 37.750 / 14.270 0.204 / 0.212 / 0.235 0.005 / 0.020 / 0.029 0.033 / 0.050 / 0.071
ZNet 120.243 / 19.860 / 9.497 0.148 / 0.150 / 0.193 0.019 / 0.054 / 0.077 0.025 / 0.054 / 0.071
AutoIV 79.648 / 20.792 / 7.362 0.253 / 0.267 / 0.273 0.000 / 0.000 / 0.000 0.032 / 0.049 / 0.091
VIV 16.644 / 6.861 / 4.575 0.033 / 0.061 / 0.117 0.033 / 0.057 / 0.106 0.040 / 0.065 / 0.087
GIV 34.207 / 11.849 / 5.966 0.141 / 0.146 / 0.169 0.017 / 0.039 / 0.058 0.028 / 0.057 / 0.087
Non-linear Mixed TrueIV 4.313 / 1.787 / 2.142 0.027 / 0.051 / 0.124 0.058 / 0.098 / 0.102 0.030 / 0.062 / 0.045
ZNet 144.440 / 18.633 / 15.304 0.167 / 0.156 / 0.189 0.036 / 0.083 / 0.107 0.201 / 0.185 / 0.221
AutoIV 94.883 / 17.895 / 17.240 0.210 / 0.222 / 0.238 0.000 / 0.000 / 0.000 0.224 / 0.210 / 0.248
VIV 15.459 / 4.695 / 3.970 0.035 / 0.058 / 0.111 0.043 / 0.068 / 0.117 0.052 / 0.062 / 0.091
GIV 56.539 / 18.722 / 9.017 0.138 / 0.145 / 0.194 0.034 / 0.043 / 0.077 0.136 / 0.126 / 0.153
Non-linear Mixed (no $U\to X$) TrueIV 16.903 / 4.576 / 3.954 0.027 / 0.051 / 0.124 0.026 / 0.057 / 0.120 0.027 / 0.055 / 0.085
ZNet 281.071 / 28.446 / 19.673 0.236 / 0.239 / 0.275 0.020 / 0.038 / 0.071 0.043 / 0.059 / 0.067
AutoIV 163.159 / 34.251 / 23.897 0.229 / 0.235 / 0.257 0.000 / 0.000 / 0.000 0.031 / 0.063 / 0.069
VIV 19.235 / 3.738 / 4.668 0.035 / 0.059 / 0.114 0.034 / 0.065 / 0.099 0.032 / 0.064 / 0.087
GIV 24.444 / 8.313 / 5.430 0.137 / 0.160 / 0.188 0.024 / 0.036 / 0.061 0.032 / 0.053 / 0.079
Non-linear No Candidate TrueIV
ZNet 54.003 / 9.565 / 4.068 0.144 / 0.186 / 0.187 0.022 / 0.061 / 0.091 0.198 / 0.195 / 0.191
AutoIV 62.921 / 16.672 / 6.355 0.202 / 0.201 / 0.236 0.000 / 0.000 / 0.000 0.322 / 0.324 / 0.318
VIV 20.382 / 4.410 / 4.932 0.034 / 0.057 / 0.115 0.031 / 0.047 / 0.096 0.040 / 0.060 / 0.090
GIV 23.479 / 3.748 / 3.176 0.145 / 0.147 / 0.182 0.029 / 0.028 / 0.069 0.091 / 0.098 / 0.099
Non-linear No Candidate (no $U\to X$) TrueIV
ZNet 100.546 / 7.451 / 6.169 0.114 / 0.121 / 0.171 0.033 / 0.063 / 0.081 0.043 / 0.052 / 0.075
AutoIV 55.722 / 7.239 / 7.235 0.215 / 0.217 / 0.246 0.000 / 0.000 / 0.000 0.033 / 0.054 / 0.073
VIV 19.903 / 7.638 / 5.995 0.033 / 0.062 / 0.118 0.039 / 0.049 / 0.108 0.039 / 0.060 / 0.073
GIV 71.468 / 20.206 / 8.957 0.151 / 0.153 / 0.190 0.021 / 0.036 / 0.066 0.041 / 0.058 / 0.086
Non-linear No Candidate (no $U$) TrueIV
ZNet 97.217 / 3.587 / 3.745 0.144 / 0.149 / 0.207 0.032 / 0.039 / 0.110 – / – / –
AutoIV 40.972 / 4.580 / 5.317 0.215 / 0.222 / 0.251 0.000 / 0.000 / 0.000 – / – / –
VIV 18.745 / 3.922 / 5.198 0.031 / 0.063 / 0.109 0.022 / 0.063 / 0.118 – / – / –
GIV 87.656 / 21.257 / 17.612 0.145 / 0.151 / 0.184 0.031 / 0.036 / 0.080 – / – / –
Figure 7: Directed acyclic graphs (DAGs) demonstrating the various data generation processes on which ZNet is evaluated. Linear and non-linear relationships are constructed for each DAG giving 18 total datasets for evaluation. Maintext results focus on cases where $U$ influences $X$ as this is more challenging, more general, and unique to ZNet.